If sec α = 5/4 and cot β = 3/4, α is acute, and 1800 < β < 2700. Then, cos (α + 2β)
1
Expert's answer
2019-07-29T12:39:05-0400

Solution

1)

"sec(\u03b1)=1\/cos(\u03b1)"

cos(α)=4/5

from the main trigonometrical identity

"sin^2(x)+cos^2(x)=1"

therefore


"sin(α)=\sqrt{1−cos^2(α)}"


"sin(α)=\sqrt{1−(4/5)^2}=|3/5|"


on a condition : α is acute, therefore

sin(α)=3/5


2)

One assumes "180^{\circ}<\beta<270^{\circ}", then cot β = 3/4.

Using the main trigonometrical identity


"cot^2(x)+1=1\/sin^2(x)"

from this expression


"sin(x)=\sqrt{1/(cot^2(x)+1)}"

"sin(β)=\sqrt{1/((3/4)^2+1)}=\sqrt{16/25}=|4/5|"

on a condition : "180^{\circ}<\beta<270^{\circ}", therefore

sin(β)=-4/5.

Similarly, as in the former case we will find cos(β):

"cos(β)=\sqrt{1−sin^2(β)}"

"cos(β)=\sqrt{1−(-4/5)^2}=|3/5|"

on a condition "180^{\circ}<\beta<270^{\circ}", therefore

cos(β)=-3/5

3)

"cos(\u03b1+2\u03b2) = cos(\u03b1)\u00b7cos(2\u03b2) - sin(\u03b1)\u00b7sin(2\u03b2)"

"cos(2\u03b2)=cos^2(\u03b2)-sin^2(\u03b2)=9\/25-16\/25=-7\/25"

"sin(2\u03b2)=2sin(\u03b2)cos(\u03b2)=2(-4\/5)(-3\/5)=24\/25"

cos(2β)=-7/25, cos(α)=4/5, sin(2β)=24/25, sin(α)=3/5


"cos(\u03b1+2\u03b2) = 4\/5\u00b7(-7\/25) - 3\/5\u00b724\/25"

"cos(\u03b1+2\u03b2) = -28\/125-72\/125=-100\/125=-4\/5"

Answer:

"cos(\u03b1+2\u03b2)=-4\/5"